OpenStudy (anonymous):

3. Use the equation for the table to predict the number of miles you could drive on 12 gallons of gas. gallons of gas used 1 2 3 4 miles driven 26 52 78 104 (Points : 1) 200 252 300 312 4. What is the slope for the following table of values? The table represents the total amount spent in a paperback book club after an introductory fee of $12. Books bought 1 3 5 7 Total Amount Spent $ 21 39 57 75 (Points : 1) 3 5 7 9

OpenStudy (anonymous):

For the first one, they are using 26 gallons of gas, so you would multiply 26(M) by 12(g). The second one I know how to make an equation if you were trying to find how many books they were buying...

OpenStudy (blurbendy):

1) 12 * 26 = 312 Each mile driven is 26 times a gallon of gas used 2) Take two "points" from the table. (1, 21) and (3, 39) for example. Use the formula to find the slope and you should get: 37 - 21 / 3 - 1 = 18 / 2 = 9 So the slope would be 9.